The leader of the association of jeepney drivers claims that the average daily take home pay for all jeepney drivers in Baguio City is 400 pesos. A random sample of 100 jeepney drivers in Baguio City was interviewed and the average daily take home pay of these drivers is found to be 425 pesos. Use a 0.05 significance level to find out if the average daily take home pay of all jeepney drivers in Baguio City is different from 400 pesos. Assume that the population variance is 92 pesos.


1
Expert's answer
2022-04-12T17:44:53-0400

"H_0: \mu=400"

"H_1: \mu \not=400"

"\sigma=\sqrt{D}=\sqrt{92}=9.59"

"z=\frac{x-\mu}{\sigma/\sqrt{n}}=\frac{425-400}{9.59/\sqrt{100}}=26.1"

p-value

"2P(Z<26.1)=2x0=0"

Since p-value < 0.05, we reject the null hypothesis. So, the average daily take home pay of all jeepney drivers in pasay city is different from php 400.